OpenStudy (anonymous):

Suppose a triangle has two sides of length 2 and 5 and that the angle between these two sides is pi/3. What is the length of the third side of the triangle?

OpenStudy (mathmale):

Surely sounds like a case of side-angle-side, doesn't it? Which trig formula would help you find the length of the third side in the quickest possible way?

OpenStudy (anonymous):

I'd assume the one that goes: a^2 = b^2 + c^2 - 2(b)(c)(cos(angle)) ?? I think that's correct...

OpenStudy (mathmale):

Dear Perfectly Capable, I'd suggest you sketch this situation for yourself, if you haven't already. yes, you're absolutely correct. Bet you can now find that "angle".

OpenStudy (mathmale):

I disagree with the "Incapable" part.

OpenStudy (anonymous):

a^2 = 2^2 +5^2 - 2*2*5*cos(pi/3) that's what I tried, but i'm getting the wrong answer. and thank you :)

OpenStudy (mathmale):

You're welcome! We want side a. We therefore apply the formula exactly as you have typewritten it: \[a ^{2}=b ^{^{2}}+c ^{2}-2bc*\cos A\] or

OpenStudy (mathmale):

\[a ^{2}=(2)^{2}+(5)^2 - 2(2)(5) (\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

a^2 = 4 + 25 - 20(cos(1.05) a^2 = 29 - 20(1) a^2 = 29 - 20 a^2 = 9 a = 3 That's how i did it (rounding the cos part of course), but it doesn't seem to be correct

OpenStudy (anonymous):

why would it be 1/2?

OpenStudy (mathmale):

Aha. It helps so much that you've shared your work. Note that angle A is pi/3, or A = 60 degrees. How do you normally find the sine or cosine of a given angle?

OpenStudy (mathmale):

Use whatever method to evaluate cos (pi/3), or cos (30 degrees).

OpenStudy (mathmale):

Very sorry, but that's cos (60 degrees).

OpenStudy (anonymous):

oh my god... i completely forgot about converting pi/3 into degrees. doing (pi/3)*(180/pi)=60 i got you now. i'm sorry... my mind is a mess

OpenStudy (anonymous):

then cos(60) = 1/2

OpenStudy (mathmale):

Note that the angle 60 degrees drawn in Q1 will have opposite side equal to Sqrt(3), adj side 1, and hyp 2. So yes, cos 60 = 1/2. You ok with that?

OpenStudy (mathmale):

With all this info, you should now be able to find a^2, and from a^2, to find a.

OpenStudy (anonymous):

a^2 = 2^2 + 5^2 - 2(2)(5)(1/2) a^2 = 4 + 25 - 10(1) a^2 = 29 - 10 a^2 = 19 a = sqrt19 which is one of the answers

OpenStudy (mathmale):

Hope you're feeling comfortable and confident regarding this work now. Happy for having had the opp to work with you.. :)

OpenStudy (anonymous):

I was trying to take the cosine of pi/3. I'll make sure to convert from now on. thank you very much, i get it :)

OpenStudy (mathmale):

My great pleasure. Bye!